OpenStudy (anonymous):

Use integration by substitution to find the area under the curve \[y={1 \over sqrt{x}+x}\] between x=1 and x=4.

OpenStudy (anonymous):

I've got \[\int_{1}^{4} {1 \over \sqrt{x}+x}\] and that's it. Got no idea for what u should be.

OpenStudy (anonymous):

forgot the dx, whatever

OpenStudy (blockcolder):

Why don't you try letting u=sqrt(x)?

OpenStudy (anonymous):

doesn't that mean that du is \[1 \over 2 \sqrt{x}\] then? I couldn't see how that could fit...

OpenStudy (blockcolder):

Well, that's correct, but you know that u^2=x so that 2u du=dx.

OpenStudy (anonymous):

oh yeah... give me a sec to think it through

OpenStudy (anonymous):

wait, what about the dx then?

OpenStudy (anonymous):

unless we use the du's sqrt{x} for the sqrt{x} in the problem and u^2 for x?

OpenStudy (blockcolder):

No, the you don't use the du's sqrt(x), but you can substitute u^2 for x.

OpenStudy (anonymous):

\[u=\sqrt{x}\] \[du=\frac{dx}{2\sqrt{x}}\] \[2\sqrt{x}du=dx\] \[2udu=dx\] \[2\int\frac{udu}{u^2+u}\]

OpenStudy (anonymous):

ahh.. I see what you did there. thanks.

OpenStudy (anonymous):

might as well change limits of integration while you are at it and get \[2\int_1^2\frac{udu}{u^2+u}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ok got it :)